# Finite-Agent Stochastic Differential Games on Large Graphs: I. The Linear-Quadratic Case

### v1

@@ 4.2.4 The construction of the Markovian NE
Using Corollary 4.10 for $Q(X)$ and combining it with the triangle inequality gives
$\text{Tr}\left[\left(Q^{\prime}(X_{2})-Q^{\prime}(X_{1})\right)^{2}\right]\leq\frac{2c^{2}}{N^{2}}\text{Tr}\left(L^{2}\left[Q(X_{2})(I+cX_{2}L)^{-1}-Q(X_{2})(I+cX_{1}L)^{-1}\right]^{2}\right)\\ +\frac{2c^{2}}{N^{2}}\text{Tr}\left(L^{2}\left[Q(X_{2})(I+cX_{1}L)^{-1}-Q(X_{1})(I+cX_{1}L)^{-1}\right]^{2}\right).$ (4.64)
@@
$\sum_{k=1}^{N}\left[Q(X_{2})\frac{1}{1+cx_{2,k}\lambda_{k}}-Q(X_{2})\frac{1}{1+cx_{1,k}\lambda_{k}}\right]^{2}$ $\leq\sum_{k=1}^{N}(1+2cx_{0})^{2}\left[\frac{c\lambda_{k}(x_{1,k}-x_{2,k})}{(1+cx_{1,k}\lambda_{k})(1+cx_{2,k}\lambda_{k})}\right]^{2}$ (4.66)
$\leq 64c^{2}(1+2cx_{0})^{2}N\|X_{2}-X_{1}\|^{2},$ (4.67)
@@
$[Q(X_{2})-Q(X_{1})]^{2}\leq\|X_{2}-X_{1}\|_{F}^{2}\ \sup_{\tiny{\mathclap{\begin{subarray}{c}-\frac{1}{4c}I\leq X\leq x_{0}I\\ X\in\mathscr{X}\end{subarray}}}}\|Q^{\prime}(X)\|_{F}^{2}\leq N^{2}\|X_{2}-X_{1}\|^{2}\ \sup_{\tiny\mathclap{\begin{subarray}{c}-\frac{1}{4c}I\leq X\leq x_{0}I\\ X\in\mathscr{X}\end{subarray}}}\|Q^{\prime}(X)\|^{2},$ (4.69)
@@
$\sup_{\tiny\mathclap{\begin{subarray}{c}-\frac{1}{4c}I\leq X\leq x_{0}I\\ X\in\mathscr{X}\end{subarray}}}\|Q^{\prime}(X)\|^{2}\leq\frac{c}{N}(1+2cx_{0})\sup_{\tiny-\frac{1}{4c}\leq x_{1},\cdots,x_{n}\leq x_{0}}\sup_{\tiny k\in[N]}\frac{\lambda_{k}}{1+cx_{k}\lambda_{k}}\leq\frac{4c}{N}(1+2cx_{0}).$ (4.70)
@@
Let $r_{1}(t),\cdots,r_{N}(t)$ denote the eigenvalues of $R(t)$, such that $r_{1}^{\prime}(t),\cdots,r_{N}^{\prime}(t)$ represent the eigenvalues of $R^{\prime}(t)$. It’s clear that $r_{1}(t),\cdots,r_{N}(t)$ are positive for $\forall t\in[0,T]$. Since any two matrices in the set $\{L,R(t),R^{\prime}(t)\}$ commute for any fixed $t\in[0,T]$, they can be simultaneously diagonalized. Rewritting equation (4.3) using the spectra of matrices, one has
@@
In what follows, we establish the concavity of $Q_{R}$, which is crucial for achieving global wellposedness. Differentiating equation (4.75) with respect to $t$ yields:
@@ 4.4 Verification of heuristic derivations
To achieve this, we propose an alternative governing system for $\eta$, showing its connection with equation (4.3), which allows for an easier verification procedure. We once again emphasize that verifying the assumptions in Remark 4.11 is unnecessary since the fact that $F_{t}^{i}$ given in (4.5) solves (2.13) is directly verified below.
@@ 4.4.1 The alternative governing equation for $\eta$
Let $R(t)$ be a solution to equation (4.3), we first check the initial conditions in equation (4.81). By Corollary 4.10, $Q^{\prime}(0)=\frac{c}{N}L$, thus $R^{\prime}(0)=\frac{1}{c}\text{Tr}\left[Q^{\prime}(0)\right]I=I$. So it remains to show
@@
with $Q_{1}$ defined in (4.81). By Corollary 4.10 and equation (4.4), one has
@@
On the other hand, a calculation of $Q_{1}^{\prime}(t)$ based on Corollary 4.10 provides
@@
$Q_{1}^{\prime}(t)=\frac{1}{c}\text{Tr}\left(Q^{\prime}(R(t))e^{-t(a+q)L}\right)\text{Tr}\left(\frac{1}{N}\left[\det\left(I+cLR(t)\right)\right]^{\frac{1}{N}}(I+cLR(t))^{-1}cLe^{-t(a+q)L}\right),$ (4.85)
@@ 4.4.2 Verification through the alternative governing equation for $\eta$
Based on equation (4.6) and Lemma 4.7(i), $P_{t}-(a+q)L$ takes values as symmetric matrices and commutes with $R_{\varphi}$ for all $\varphi\in\text{Aut}(G)$ for any $t\in[0,T]$. Lemma 4.7(ii) implies that $e_{k}^{\operatorname{T}}[P_{t}-(a+q)L]e_{k}=\tau(t)$ for each $k\in[N]$. Combining this with equation (4.5) yields
@@
This affirms the validity of equation (4.20) within the context of verification, which has been a key element throughout the heuristic construction of the equilibrium.
@@
$-\frac{\tau^{\prime}(t)}{\tau(t)}[P_{t}-(a+q)L]e_{i}e_{i}^{\operatorname{T}}[P_{t}-(a+q)L]+\dot{P}_{t}e_{i}e_{i}^{T}[P_{t}-(a+q)L]+[P_{t}-(a+q)L]e_{i}e_{i}^{\operatorname{T}}\dot{P}_{t}\\ -P_{t}[P_{t}-(a+q)L]e_{i}e_{i}^{\operatorname{T}}[P_{t}-(a+q)L]-[P_{t}-(a+q)L]e_{i}e_{i}^{\operatorname{T}}[P_{t}-(a+q)L]P_{t}\\ +\tau(t)[P_{t}-(a+q)L]e_{i}e_{i}^{\operatorname{T}}[P_{t}-(a+q)L]=0.$ (4.88)
Since any two matrices in the set $\{P_{t},L,R(T-t),R^{\prime}(T-t)\}$ commute for any fixed $t\in[0,T]$, they are simultaneously diagonalizable. Recall that we denote $\lambda_{1},\ldots,\lambda_{N}$ as the eigenvalues of $L$, and $\rho^{1}_{t},\ldots,\rho^{N}_{t}$ as the eigenvalues of $P_{t}$. According to equation (4.81), the $j$-th eigenvalues of $R(T-t)$ and $R^{\prime}(T-t)$ are $\int_{t}^{T}\sqrt{\frac{Q_{1}^{\prime}(T-s)}{c}}e^{-(T-s)(a+q)\lambda_{j}}\,\mathrm{d}s$ and $\sqrt{\frac{Q_{1}^{\prime}(T-t)}{c}}e^{-(T-t)(a+q)\lambda_{j}}$ respectively. As a result, equation (4.6) can be expressed in the spectral form as
@@
$-\frac{\tau^{\prime}(t)}{\tau(t)}[\rho^{j}_{t}-(a+q)\lambda_{j}][\rho^{k}_{t}-(a+q)\lambda_{k}]+\dot{\rho}^{j}_{t}[\rho^{k}_{t}-(a+q)\lambda_{k}]+[\rho^{j}_{t}-(a+q)\lambda_{j}]\dot{\rho}^{k}_{t}\\ -\rho^{j}_{t}[\rho^{j}_{t}-(a+q)\lambda_{j}][\rho^{k}_{t}-(a+q)\lambda_{k}]-[\rho^{j}_{t}-(a+q)\lambda_{j}][\rho^{k}_{t}-(a+q)\lambda_{k}]\rho^{k}_{t}\\ +\tau(t)[\rho^{j}_{t}-(a+q)\lambda_{j}][\rho^{k}_{t}-(a+q)\lambda_{k}]=0,\ \forall j,k\in[N],$ (4.90)
@@
which corresponds to the $k$-th eigenvalue of the matrix $-R^{\prime\prime}(T-t)[R^{\prime}(T-t)]^{-1}-(a+q)L$, using equation (4.81).
@@
$\tau(t)=-\frac{\mathrm{d}}{\mathrm{d}t}\log\left[\prod_{k=1}^{N}\left(1+\lambda_{k}\int_{t}^{T}g_{k}(s)\,\mathrm{d}s\right)\right]^{\frac{1}{N}}=-\frac{\mathrm{d}}{\mathrm{d}t}\log\left[\det\left(I+cL\int_{t}^{T}R^{\prime}(T-s)\,\mathrm{d}s\right)\right]^{\frac{1}{N}}.$ (4.97)
@@
$-R^{\prime\prime}(T-t)[R^{\prime}(T-t)]^{-1}-(a+q)L=-\frac{Q_{1}^{\prime\prime}(T-t)}{2Q_{1}^{\prime}(T-t)}I.$ (4.101)
@@ 5 Numerical Experiments
In this section, we conduct several numerical experiments concerning the NE of the game introduced in Section 2.1. Section 5.1 computes the semi-explicit equilibrium derived in Theorem 4.7 under various vertex-transitive graphs, and numerically verifies its consistency with solving the Riccati system (2.13). Implied by Corollary 3.8, fictitious play is shown to be a generally applicable technique for any connected graphs with provable convergence. In Section 5.2, we visualize this convergence procedure and quantify its rate of convergence for different graphs. Lastly, in Section 5.3, a comparison of the time complexity and the running time among previously mentioned numerical methods is presented.
@@
[…]
@@
These graphs are of great research interest in their own, and we refer readers to [Chu97, GR01] for their definitions, constructions, and properties.
@@
MAE $3.88\times 10^{-6}$ $2.09\times 10^{-5}$ $7.12\times 10^{-6}$ $4.67\times 10^{-6}$ $1.92\times 10^{-6}$ $3.16\times 10^{-5}$
MRE $5.07\times 10^{-6}$ $8.96\times 10^{-5}$ $1.44\times 10^{-5}$ $8.21\times 10^{-4}$ $5.30\times 10^{-6}$ $1.16\times 10^{-3}$
@@
[…]
@@
Figure 2: Comparisons of equilibrium state (left panel) and strategy (right panel) trajectories for $N=50$ players in the linear-quadratic game on the cycle graph $G=C_{50}$. In both panels, the colored solid lines represent the baseline solution (by numerically solving the Riccati system (2.13)), and the colored circles are obtained by numerically solving the semi-explicit solution in Theorem 4.7. For the sake of clarity, only trajectories of four randomly selected players are plotted.
@@ 5.2 Numerical convergence of fictitious play
[…]
$K_{m,m}:\text{the $2m$-vertex complete bipartite graph with partitions of size $m$ and $m$},$ (5.11)
$\text{RSG}^{s}_{N}:\text{the $N$-vertex random sparse graph with edge density $s$},$ (5.12)
[…]
@@
MAE $2.19\times 10^{-3}$ $8.58\times 10^{-2}$ $1.50\times 10^{-1}$ $1.76\times 10^{-2}$
MRE $3.21\times 10^{-3}$ $1.02\times 10^{-1}$ $1.69\times 10^{-1}$ $2.26\times 10^{-2}$
@@
MAE $5.62\times 10^{-2}$ $4.60\times 10^{-2}$ $4.58\times 10^{-2}$ $3.69\times 10^{-2}$ $3.70\times 10^{-2}$ $3.68\times 10^{-2}$
MRE $7.23\times 10^{-2}$ $6.03\times 10^{-2}$ $5.98\times 10^{-2}$ $4.82\times 10^{-2}$ $4.82\times 10^{-2}$ $4.83\times 10^{-2}$
The MAEs and MREs shown in Tables 23 are consistently small, implying the numerical convergence of FP, even when the model parameters are deliberately chosen to not satisfy condition (3.30) of Corollary 3.8. We observe that the magnitude of errors varies for different graph structures. To better compare the convergence rates, we plot $\log\text{MAE}(\check{F}^{\cdot,k-1};\check{F}^{\cdot,k})$ vs. the FP stage index $k$, as shown in Figure 3. The log-MAE curves are nearly straight lines with slopes depending on the graph structures. This implies a contraction-mapping-like structure
$\text{MAE}(\check{F}^{\cdot,k};\check{F}^{\cdot,k+1})\approx C_{G}\cdot\text{MAE}(\check{F}^{\cdot,k-1};\check{F}^{\cdot,k}),\ \forall k\in\mathbb{N},$ (5.14)
where $C_{G}\in(0,1)$ is a constant that depends on the graph $G$. A simple linear regression gives
$C_{K_{50}}=-4.55,\;C_{C_{50}}=-1.94,\;C_{S_{50}}=-2.01,\;C_{K_{25,25}}=-2.61,\;C_{\text{RSG}^{0.125}_{50}}=-2.06,\;C_{\text{RMJ}^{6}_{50}}=-2.07.$
@@
Figure 3: Comparisons of log-MAE curves for the linear-quadratic game with $N=50$ players on different graphs. The behavior at the tail of the log-MAE curves for the complete graphs is due to the numerical round-off error.

### v2

@@ 4.1.2 Proof roadmap
Using Corollary A.5 for $Q(X)$ and combining it with the triangle inequality gives
$\hskip-7.0pt\text{Tr}\left[\left(Q^{\prime}(X_{2})-Q^{\prime}(X_{1})\right)^{2}\right]\leq\frac{2c^{2}}{N^{2}}\text{Tr}\left(L^{2}\left[Q(X_{2})(I+cX_{2}L)^{-1}\hskip-7.0pt-Q(X_{2})(I+cX_{1}L)^{-1}\right]^{2}\right)\\ +\frac{2c^{2}}{N^{2}}\text{Tr}\left(L^{2}\left[Q(X_{2})(I+cX_{1}L)^{-1}-Q(X_{1})(I+cX_{1}L)^{-1}\right]^{2}\right).$ (4.26)
@@
$\sum_{k=1}^{N}$ $\left[Q(X_{2})\frac{1}{1+cx_{2,k}\lambda_{k}}-Q(X_{2})\frac{1}{1+cx_{1,k}\lambda_{k}}\right]^{2}$ (4.28)
$\leq\sum_{k=1}^{N}(1+2cx_{0})^{2}\left[\frac{c\lambda_{k}(x_{1,k}-x_{2,k})}{(1+cx_{1,k}\lambda_{k})(1+cx_{2,k}\lambda_{k})}\right]^{2}$ (4.29)
$\leq 64c^{2}(1+2cx_{0})^{2}N\|X_{2}-X_{1}\|^{2}.$ (4.30)
@@
$Q(X_{2})-Q(X_{1})]^{2}\leq\|X_{2}-X_{1}\|_{F}^{2}\ \sup_{\tiny{\mathclap{\begin{subarray}{c}0\leq X\leq x_{0}I\\ X\in\mathscr{X}\end{subarray}}}}\|Q^{\prime}(X)\|_{F}^{2}\leq N^{2}\|X_{2}-X_{1}\|^{2}\ \sup_{\tiny\mathclap{\begin{subarray}{c}0\leq X\leq x_{0}I\\ X\in\mathscr{X}\end{subarray}}}\|Q^{\prime}(X)\|^{2},$ (4.32)
@@
$\sup_{\tiny\mathclap{\begin{subarray}{c}0\leq X\leq x_{0}I\\ X\in\mathscr{X}\end{subarray}}}\|Q^{\prime}(X)\|^{2}\leq\frac{c}{N}(1+2cx_{0})\sup_{\tiny 0\leq x_{1},\cdots,x_{n}\leq x_{0}}\sup_{\tiny k\in[N]}\frac{\lambda_{k}}{1+cx_{k}\lambda_{k}}\leq\frac{4c}{N}(1+2cx_{0}).$ (4.33)
@@
Let $r_{1}(t),\cdots,r_{N}(t)$ denote the eigenvalues of $R(t)$, such that $r_{1}^{\prime}(t),\cdots,r_{N}^{\prime}(t)$ represent the eigenvalues of $R^{\prime}(t)$. It’s clear that $r_{1}(t),\cdots,r_{N}(t)$ are positive for $\forall t\in[0,T]$. Since any two matrices in the set $\{L,R(t),R^{\prime}(t)\}$ commute for any fixed $t\in[0,T]$, they can be simultaneously diagonalized. Rewriting equation (4.3) using the spectra of matrices, one has
@@
In what follows, we establish the concavity of $Q_{R}$, which is crucial for achieving global well-posedness. Differentiating equation (4.38) with respect to $t$ yields:
@@ 4.3 Verification
To achieve this, we propose an alternative governing system for $\eta$, showing its connection with equation (4.3), which allows for an easier verification procedure. We once again emphasize that verifying the assumptions in Remark A.6 is unnecessary since the fact that $F_{t}^{i}$ given in (4.5) solves (2.14) is directly verified below.
@@ 4.3.1 The alternative governing equation for $\eta$
Let $R(t)$ be a solution to equation (4.3), we first check the initial conditions in equation (4.44). By Corollary A.5, $Q^{\prime}(0)=\frac{c}{N}L$, thus $R^{\prime}(0)=\frac{1}{c}\text{Tr}\left[Q^{\prime}(0)\right]I=I$. So it remains to show
@@
with $Q_{1}$ defined in (4.44). By Corollary A.5 and equation (4.4), one has
@@
On the other hand, a calculation of $Q_{1}^{\prime}(t)$ based on Corollary A.5 provides
@@
$Q_{1}^{\prime}(t)=\frac{1}{c}\text{Tr}\left(Q^{\prime}(R(t))e^{-t(a+q)L}\right)\\ \text{Tr}\left(\frac{1}{N}\left[\det\left(I+cLR(t)\right)\right]^{\frac{1}{N}}(I+cLR(t))^{-1}cLe^{-t(a+q)L}\right),$ (4.48)
@@ 4.3.2 Verification through the alternative governing equation for $\eta$
Based on equation (4.6) and Lemma A.2(i), $P_{t}-(a+q)L$ takes values as symmetric matrices and commutes with $R_{\varphi}$ for all $\varphi\in\text{Aut}(G)$ for any $t\in[0,T]$. Lemma A.2(ii) implies that $e_{k}^{\operatorname{T}}[P_{t}-(a+q)L]e_{k}=\tau(t)$ for each $k\in[N]$. Combining this with equation (4.5) yields
@@
This affirms the validity of equation (A.1) within the context of verification, which has been a key element throughout the heuristic construction of the equilibrium.
@@
$\frac{\tau^{\prime}(t)}{\tau(t)}[P_{t}-(a+q)L]e_{i}e_{i}^{\operatorname{T}}[P_{t}-(a+q)L]+\dot{P}_{t}e_{i}e_{i}^{T}[P_{t}-(a+q)L]+[P_{t}-(a+q)L]e_{i}e_{i}^{\operatorname{T}}\dot{P}_{t}\\ -P_{t}[P_{t}-(a+q)L]e_{i}e_{i}^{\operatorname{T}}[P_{t}-(a+q)L]-[P_{t}-(a+q)L]e_{i}e_{i}^{\operatorname{T}}[P_{t}-(a+q)L]P_{t}\\ +\tau(t)[P_{t}-(a+q)L]e_{i}e_{i}^{\operatorname{T}}[P_{t}-(a+q)L]=0.$ (4.51)
Since any two matrices in the set $\{P_{t},L,R(T-t),R^{\prime}(T-t)\}$ commute for any fixed $t\in[0,T]$, they are simultaneously diagonalizable. Recall that we denote $\lambda_{1},\ldots,\lambda_{N}$ as the eigenvalues of $L$, and $\rho^{1}_{t},\ldots,\rho^{N}_{t}$ as the eigenvalues of $P_{t}$. According to equation (4.44), the $j$-th eigenvalues of R(Tt)R(T-t) and $R^{\prime}(T-t)$ are $\int_{t}^{T}\sqrt{\frac{Q_{1}^{\prime}(T-s)}{c}}e^{-(T-s)(a+q)\lambda_{j}}\,\mathrm{d}s$ and $\sqrt{\frac{Q_{1}^{\prime}(T-t)}{c}}e^{-(T-t)(a+q)\lambda_{j}}$ respectively. As a result, equation (4.6) can be expressed in the spectral form as
@@
$\frac{\tau^{\prime}(t)}{\tau(t)}[\rho^{j}_{t}-(a+q)\lambda_{j}][\rho^{k}_{t}-(a+q)\lambda_{k}]+\dot{\rho}^{j}_{t}[\rho^{k}_{t}-(a+q)\lambda_{k}]+[\rho^{j}_{t}-(a+q)\lambda_{j}]\dot{\rho}^{k}_{t}\\ -\rho^{j}_{t}[\rho^{j}_{t}-(a+q)\lambda_{j}][\rho^{k}_{t}-(a+q)\lambda_{k}]-[\rho^{j}_{t}-(a+q)\lambda_{j}][\rho^{k}_{t}-(a+q)\lambda_{k}]\rho^{k}_{t}\\ +\tau(t)[\rho^{j}_{t}-(a+q)\lambda_{j}][\rho^{k}_{t}-(a+q)\lambda_{k}]=0,\ \forall j,k\in[N],$ (4.53)
@@
which corresponds to the $k$-th eigenvalue of the matrix $R^{\prime\prime}(T-t)[R^{\prime}(T-t)]^{-1}-(a+q)L$, using equation (4.44).
@@
$\tau(t)=-\frac{\mathrm{d}}{\mathrm{d}t}\log\left[\prod_{k=1}^{N}\left(1+\lambda_{k}\int_{t}^{T}g_{k}(s)\,\mathrm{d}s\right)\right]^{\frac{1}{N}}\\ =-\frac{\mathrm{d}}{\mathrm{d}t}\log\left[\det\left(I+cL\int_{t}^{T}R^{\prime}(T-s)\,\mathrm{d}s\right)\right]^{\frac{1}{N}}.$ (4.60)
@@
$R^{\prime\prime}(T-t)[R^{\prime}(T-t)]^{-1}-(a+q)L=-\frac{Q_{1}^{\prime\prime}(T-t)}{2Q_{1}^{\prime}(T-t)}I.$ (4.64)
@@ 5 Numerical Experiments
In this section, we conduct several numerical experiments concerning the NE of the game introduced in Section 2.1. Section 5.1 computes the semi-explicit equilibrium derived in Theorem 4.2 under various vertex-transitive graphs, and numerically verifies its consistency with solving the Riccati system (2.14). Implied by Corollary 3.8, fictitious play is shown to be a generally applicable technique for any connected graphs with provable convergence. In Section 5.2, we visualize this convergence procedure and quantify its convergence rate for different graphs. Lastly, in Section 5.3, a comparison of the time complexity and the running time among previously mentioned numerical methods is presented.
@@
where the initial state $x_{0}$ is randomly sampled from a uniform distribution $x_{0}^{i}\overset{\text{i.i.d.}}{\sim}\mathcal{U}(-1,1),\ \forall i\in[N]$. The remaining model parameters $(q,\varepsilon)$, $N$, $T$ and $G$ vary and will be specified in Table 1. Throughout the experiments, the relationship $q^{2}=\varepsilon$ is maintained and $G$ is taken as one of the following vertex-transitive graphs:
$KG_{5,2}:\text{the $10$-vertex Petersen graph},\quad K_{N}:\text{the $N$-vertex complete graph},$ (5.2)
$C_{N}:\text{the $N$-vertex cycle graph},\qquad\qquad Q_{k}:\text{the $2^{k}$-vertex hypercube graph},$ (5.3)
@@
These graphs are of great research interest on their own, and we refer readers to [Chu97, GR01] for their definitions, constructions, and properties.
@@
MAE 3.88e63.88e{-6} 2.09e52.09e{-5} 7.12e67.12e{-6} 4.67e64.67e{-6} 1.92e61.92e{-6} 3.16e53.16e{-5}
MRE 5.07e65.07e{-6} 8.96e58.96e{-5} 1.44e51.44e{-5} 8.21e48.21e{-4} 5.30e65.30e{-6} 1.16e31.16e{-3}
@@
$\hat{X}^{i}_{t+h}=\hat{X}^{i}_{t}+\left[a\left(\frac{1}{\sqrt{d_{v_{i}}}}\sum_{j:v_{j}\sim v_{i}}\frac{1}{\sqrt{d_{v_{j}}}}\hat{X}^{j}_{t}-\hat{X}^{i}_{t}\right)+\hat{\alpha}^{i}_{t}\right]h+\sigma\sqrt{h}\xi^{i}_{t},\ \xi^{i}_{t}\overset{\text{i.i.d.}}{\sim}\mathcal{N}(0,1),$ (5.6)
@@
Figure 2: Comparisons of equilibrium state (left panel) and strategy (right panel) trajectories for $N=50$ players in the linear-quadratic game on the cycle graph $G=C_{50}$. In both panels, the colored solid lines represent the baseline solution (by numerically solving the Riccati system (2.14)), and the colored circles are obtained by numerically solving the semi-explicit solution in Theorem 4.2. For the sake of clarity, only trajectories of four randomly selected players are plotted
@@ 5.2 Numerical convergence of fictitious play
where the initial state $x_{0}$ is randomly sampled from a uniform distribution $x_{0}^{i}\overset{\text{i.i.d.}}{\sim}\mathcal{U}(-1,1),\ \forall i\in[N]$. Throughout the experiments, $G$ is taken as one of the following graphs, with the specific choice given in Tables 23:
$K_{N}:\text{the $N$-vertex complete graph},\qquad C_{N}:\text{the $N$-vertex cycle graph},$ (5.9)
$S_{N}:\text{the $N$-vertex star graph},\qquad\qquad$ (5.10)
@@
$\text{RMJ}^{d}_{N}:\text{the $N$-vertex $d$-regular random Ramanujan graph.}$ (5.13)
@@
MAE 2.19e32.19e{-3} 8.58e28.58e{-2} 1.50e11.50e{-1} 1.76e21.76e{-2}
MRE 3.21e33.21e{-3} 1.02e11.02e{-1} 1.69e11.69e{-1} 2.26e22.26e{-2}
@@
MAE 5.62e25.62e{-2} 4.60e24.60e{-2} 4.58e24.58e{-2} 3.69e23.69e{-2} 3.70e23.70e{-2} 3.68e23.68e{-2}
MRE 7.23e27.23e{-2} 6.03e26.03e{-2} 5.98e25.98e{-2} 4.82e24.82e{-2} 4.82e24.82e{-2} 4.83e24.83e{-2}
The MAEs and MREs shown in Tables 23 are consistently small, implying the numerical convergence of FP, even when the model parameters are deliberately chosen not to satisfy condition (3.31) of Corollary 3.8. We observe that the magnitude of errors varies for different graph structures. To better compare the convergence rates, we plot $\log\text{MAE}(\check{F}^{\cdot,k-1};\check{F}^{\cdot,k})$ vs. the FP stage index $k$, as shown in Figure 3. The log-MAE curves are nearly straight lines with slopes depending on the graph structures. This implies a contraction-mapping-like structure
$\text{MAE}(\check{F}^{\cdot,k};\check{F}^{\cdot,k+1})\approx e^{C_{G}}\cdot\text{MAE}(\check{F}^{\cdot,k-1};\check{F}^{\cdot,k}),\ \forall k\in\mathbb{N},$ (5.14)
where $C_{G}\in(-\infty,0)$ is a constant that depends on the graph $G$. A simple linear regression gives
$C_{K_{50}}=-4.55,\;$ $C_{C_{50}}=-1.94,\;$ $C_{S_{50}}=-2.01,\;$
$C_{K_{25,25}}=-2.61,\;$ $C_{\text{RSG}^{0.125}_{50}}=-2.06,\;$ $C_{\text{RMJ}^{6}_{50}}=-2.07.$
@@
Figure 3: Comparisons of log-MAE curves for the linear-quadratic game with $N=50$ players on different graphs. The behavior at the tail of the log-MAE curves for the complete graphs is due to the numerical round-off error